OpenStudy (anonymous):

the legs of a right triangle are 15cm. and 36cm. what is the length of the hypotenuse? a. 21 b. 39 c. 51 d. 57

OpenStudy (solomonzelman):

\(\huge\color{black}{ \sf a^2+b^2=c^2~~-> }\) \(\huge\color{blue}{ \sf 15^2+36^2=c^2 }\) \(\huge\color{blue}{ \sf 15^2+36^2=c^2 }\) \(\huge\color{blue}{ \sf 225+1296=c^2 }\) \(\huge\color{blue}{ \sf 1521=c^2 }\) \(\huge\color{blue}{ \sf ±39=c }\) \(\color{black}{ \sf distance~~~~can't~~~~be~~~~negative,~~~~it~~~~can~~~~only~~~~be~~~~positive.~~~~So, }\) \(\color{black}{ \sf your~~~~answer~~~~would~~~~be }\) \(\color{black}{ 39 }\), \(\color{black}{ \sf which~~~~is~~~~option }\) \(\color{black}{ B }\) \(\color{black}{ \sf in~~~your~~choices.}\)

OpenStudy (anonymous):

Thanks you were right!!

OpenStudy (solomonzelman):

Yeah, I bet I was:) xD come here if you need more help.

OpenStudy (anonymous):

k! i probably will cause i'm homeschooled

OpenStudy (solomonzelman):

Home-schooled? Alrighty... , it must be difficult there without a teacher near. I can imagine you having to wait a day or 2 (times even 3 or 4) for a teacher to web-reply. Good luck with school !

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

but i've been homeschooled my whole life so normally i don't have problems like that!

OpenStudy (anonymous):

but i still hate math teaching it to ur self is even harder